# DivShift: Exploring Domain-Specific Distribution Shift in Volunteer-Collected Biodiversity Datasets

### v1

@@ 1 Introduction
The world’s biodiversity is under threat from climate and land-use change [50, 46]. Biodiversity helps ecosystems combat climate change by improving carbon sequestration [70] and novel climate adaptation [56]. Therefore, monitoring the world’s biodiversity via automated tools [53] is critical to mitigate climate change’s effects on the natural world. Building machine learning tools for this automated monitoring requires large volumes of natural world imagery [28]. Participatory science applications—where users can upload and identify photos of species in their natural environments—have surged in popularity [21, 2, 23] (see Appdx. 1.1 for participatory science overview). These applications now provide sufficient finely-labeled imagery data to build large-scale biodiversity datasets [65, 31, 63, 58, 26, 29] and train deep learning models for automated biodiversity monitoring tasks such as species recognition [65, 26, 29, 31], species distribution modeling [63, 31, 58, 28], novel species identification [60], and visual question answering [65].
@@ 2 Related Works
Biases in volunteer-collected biodiversity datasets. Collection of large-scale volunteer datasets are subject to social and ecological filters [11, 33], which inject many types of bias into biodiversity datasets built from these collections [11, 22, 34] (see Appdx. 1.2 for further definition). In this work, we focus on four kinds of bias common to volunteer-collected biodiversity datasets: spatial, temporal, observation quality, and sociopolitical (Fig. 1). Spatial bias includes observer preferences to sampling easy-to-access greenspaces in urban areas [30, 42, 2, 22, 23, 11] (see Appdx. 1.3.1). Temporal bias [2, 23, 62, 57, 17, 32] includes a skew towards more observations on weekends when observers are free from work [22, 16, 15] (see Appdx. 1.3.2). Observation quality bias [23, 64, 43, 6] manifests as a small but dedicated group of users that tend to observe more species in more diverse habitats [22, 55] (see Appdx. 1.3.3). Lastly, sociopolitical bias in who has access to the resources, time, and areas to collect biodiversity observations [5, 24, 41, 19, 12, 11, 9, 14, 59, 48] includes a skew towards whiter, wealthier, and older observers [40, 51] (see Appdx. 1.3.4).
@@ 3 DivShift Framework and DivShift-NAWC Dataset
We focus on spatial, temporal, observation quality, and sociopolitical bias (Fig. 1). For spatial bias, we use the Human Footprint Index (HFI) [44] where observations in wilderness are HFI <= 1 and modified are HFI >= 4 (Appdx. 3.1.1). For temporal bias we compare observations from the City Nature Challenge (CNC) [21, 32] to those not (Appdx. 3.1.2). For observation quality bias, compare engaged observers ($>$ 1,000 observations) to casual observers ($<$ 50 observations) [22] (Appdx. 3.1.3). Finally for sociopolitical bias, we compare two high-resource states ($>1$ mil. images, California-C) and British Columbia-BC) to two low-resource states ($<50,000$ images, Sonora-SO and Yukon-YT) [54] (Appdx. 3.1.4). We also report accuracies on baseline train / test partitions as commonly executed in other natural world imagery datasets [65, 13, 31, 20] (Appdx. 3.1.5).
@@
Model Training and Evaluation Metrics: To measure the underlying data partition distribution shifts, we filter each bias partition (e.g. modified cities vs. wilderness) to only species shared between the partitions, and calculated the Jensen-Shannon Distance (JSD) between the label distributions (Appdx. 3.2). For each bias partition, we split images into either train (80%) or test (20%) unless otherwise specified, and train a ResNet18 for 10 epochs on a given train partition (see Appdx. 4 for training details). For each model, we report Top1 accuracy per-observation (Top1-Obs), Top1 accuracy per-species (Top1-Spec), a new Top1 accuracy metric weighted by rarity (Top1-Wgt), and Top1 accuracy by land use category (Top1-LUC) (see Appdx. 5 for metric definitions).
@@ 4 Results and Discussion
For all bias partitions, surprisingly all biases are weak, with performance drop greater than the underlying JSD of the partition (Tables A3, 1(b)). For the spatial partition, there is a larger distribution shift from wild to modified observations than vice-versa (Table 1(b)). However, training on observations from less-disturbed habitats leads to worse performance than training in areas of high human activity(Tables A3, 1(a)). For the temporal partition, the distribution shift is largely symmetric (Table 1(b)), but training on City Nature leads universally worse performance on observations from outside the Challenge, while training with observations from outside the Challenge leads to better performance except when accounting for land use (Tables A3, 1(a)). For the observation quality partition we see that the distribution shift between both the engaged and casual observer partitions are surprisingly symmetric (Table 1(b)). However, the model trained using images from engaged observers showed increased accuracy across the board, while the model trained on casual observations saw a universal decrease (Tables A3, 1(a)). For the sociopolitical bias, the distribution shift is greater for CA to SO than BC to YT (Table 1(b)), yet surprisingly the decrease in performance is less pronounced for CA to SO than BC to YT, except when correcting for land use type (Tables A3, 1(a)). Across the four splits, model accuracies (Table A3) are well within the range of the baseline splits (Table A2), implying that the bias splits are not dramatically skewed compared to previous approaches.
@@
Table 1: (1(a)) Performance differences for all partitions (absolute values in Table A3). Blue values indicate positive bias, red values indicate negative bias. Obs = Observation, Spec=Species, Wgt=Weighted, LUC=Land Use Category, CNC = City Nature Challenge, diff = difference. (1(b)) Jenson-Shannon differences between partitions. JSD = Jensen-Shannon Difference.
Climate Change Impact: Our findings on the DivShift-NAWC dataset suggest five recommendations for training deep learning models on voluminous but noisy volunteer-collected biodiversity datasets. First, observations from urban areas provide useful training signal even for wild areas for species found in both kinds of habitats. Second, using many noisy observations with a random sampling pattern is better than using fewer observations with a more structured sampling pattern. Third, using data from more engaged observers is generally better than less-engaged ones. Fourth, even though sociopolitical boundaries can have a significant effect on performance, training on diverse and extremely well-sampled high-resource regions like CA can still provide some predictive benefit for low-resource regions even if geographically far away.
@@
We first and foremost thank all of the participants who volunteered and contributed observations on iNaturalist. We further thank Noah Goodman and Gabriel Poesia for their comments and discussion, and we further thank CoCoLab for donating compute for this project. We also thank the TomKat Center for Sustainable Energy, The Fulbright Brasil Commission, Carnegie Science, the Howard Hughes Medical Institute, and the University of California, Berkeley for funding support for this research. This research was also funded by the NSF Graduate Research Fellowship DGE-1656518 (L.G.), the TomKat Graduate Fellowship for Translational Research (L.G.), the Krupp Internship Program for Stanford Students in Germany (E.S.), and T.K. and S.S. acknowledge funding from the German Research Foundation (DFG) within the framework of BigPlantSens (Assessing the Synergies of Big Data and Deep Learning for the Remote Sensing of Plant Species; project no. 444524904) and PANOPS (Revealing Earth’s plant functional diversity with citizen science; project no. 504978936). This research was further funded by the Office of the Director of the National Institutes of Health’s Early Investigator Award (1DP5OD029506-01, M.E.-A.). Compute for this project was performed on the Calc cluster at Carnegie Science and the Stanford SC Compute Cluster.

### v2

@@ 1 Introduction
The world’s biodiversity is under threat from climate and land-use change [96, 90]. Biodiversity helps ecosystems combat climate change by improving carbon sequestration [137] and novel climate adaptation [105]. Therefore, monitoring the world’s biodiversity via automated tools [100] is critical to mitigate climate change’s effects on the natural world. Building machine learning tools for this automated monitoring requires large volumes of natural world imagery [53]. Participatory science applications—where users can upload and identify photos of species in their natural environments—have surged in popularity [37, 3, 39] (see Appdx. 5.1.1 for participatory science overview). These applications now provide sufficient finely-labeled imagery data to build large-scale biodiversity datasets [125, 60, 122, 107, 50, 55] and train deep learning models for automated biodiversity monitoring tasks such as species recognition [125, 50, 55, 60], species distribution modeling [122, 60, 107, 53], novel species identification [115], and visual question answering [125].
@@ 2 Related Works
Biases in volunteer-collected biodiversity datasets. Collection of large-scale volunteer datasets are subject to social and ecological filters [21, 63], which inject many types of bias into biodiversity datasets built from these collections [21, 38, 64] (see Appdx. 5.1.2 for further definition). In this work, we focus on four kinds of bias common to volunteer-collected biodiversity datasets: spatial, temporal, observation quality, and sociopolitical (Fig. 1). Spatial bias includes observer preferences to sampling easy-to-access greenspaces in urban areas [56, 83, 3, 38, 39, 21] (see Appdx. 5.1.3). Temporal bias [3, 39, 119, 106, 30, 61] includes a skew towards more observations on weekends when observers are free from work [38, 29, 27] (see Appdx. 5.1.3). Observation quality bias [39, 124, 85, 10] manifests as a small but dedicated group of users that tend to observe more species in more diverse habitats [38, 104] (see Appdx. 5.1.3). Lastly, sociopolitical bias in who has access to the resources, time, and areas to collect biodiversity observations [9, 41, 80, 34, 24, 21, 16, 26, 111, 93] includes a skew towards whiter, wealthier, and older observers [79, 97] (see Appdx. 5.1.3).
@@ 3 DivShift Framework and DivShift-NAWC Dataset
We focus on spatial, temporal, observation quality, and sociopolitical bias (Fig. 1). For spatial bias, we use the Human Footprint Index (HFI) [87] where observations in wilderness are HFI <= 1 and modified are HFI >= 4 (Appdx. 5.3.2). For temporal bias we compare observations from the City Nature Challenge (CNC) [37, 61] to those not (Appdx. 5.3.3). For observation quality bias, compare engaged observers ($>$ 1,000 observations) to casual observers ($<$ 50 observations) [38] (Appdx. 5.3.4). Finally for sociopolitical bias, we compare two high-resource states ($>1$ mil. images, California-C) and British Columbia-BC) to two low-resource states ($<50,000$ images, Sonora-SO and Yukon-YT) [102] (Appdx. 5.3.5). We also report accuracies on baseline train / test partitions as commonly executed in other natural world imagery datasets [125, 25, 60, 35] (Appdx. 5.3.6).
@@
Model Training and Evaluation Metrics: To measure the underlying data partition distribution shifts, we filter each bias partition (e.g. modified cities vs. wilderness) to only species shared between the partitions, and calculated the Jensen-Shannon Distance (JSD) between the label distributions (Appdx. 5.5). For each bias partition, we split images into either train (80%) or test (20%) unless otherwise specified, and train a ResNet18 for 10 epochs on a given train partition (see Appdx. 5.6 for training details). For each model, we report Top1 accuracy per-observation (Top1-Obs), Top1 accuracy per-species (Top1-Spec), a new Top1 accuracy metric weighted by rarity (Top1-Wgt), and Top1 accuracy by land use category (Top1-LUC) (see Appdx. 5.7 for metric definitions).
@@ 4 Results and Discussion
For all bias partitions, surprisingly all biases are weak, with performance drop greater than the underlying JSD of the partition (Tables 1(a), 1(b)). For the spatial partition, there is a larger distribution shift from wild to modified observations than vice-versa (Table 1(b)). However, training on observations from less-disturbed habitats leads to worse performance than training in areas of high human activity (Table 1(a)). For the temporal partition, the distribution shift is largely symmetric (Table 1(b)), but training on City Nature leads universally worse performance on observations from outside the Challenge, while training with observations from outside the Challenge leads to better performance except when accounting for land use (Table 1(a)). For the observation quality partition we see that the distribution shift between both the engaged and casual observer partitions are surprisingly symmetric (Table 1(b)). However, the model trained using images from engaged observers showed increased accuracy across the board, while the model trained on casual observations saw a universal decrease (Table 1(a)). For the sociopolitical bias, the distribution shift is greater for CA to SO than BC to YT (Table 1(b)), yet surprisingly the decrease in performance is less pronounced for CA to SO than BC to YT, except when correcting for land use type (Table 1(a)). Across the four splits, model accuracies (Table A3) are well within the range of the baseline splits (Table A2), implying that the bias splits are not dramatically skewed compared to previous approaches.
@@
Table 1: (1(a)) Performance differences for all partitions (absolute values in Table A3). Blue values indicate positive bias, red values indicate negative bias. Obs = Observation, Spec=Species, Wgt=Weighted, LUC=Land Use Category, CNC = City Nature Challenge, diff = difference. (1(b)) Jensen-Shannon differences between partitions. JSD = Jensen-Shannon Difference.
Climate Change Impact: Our findings on the DivShift-NAWC dataset suggest four recommendations for training deep learning models on voluminous but noisy volunteer-collected biodiversity datasets. First, observations from urban areas provide useful training signal even for wild areas for species found in both kinds of habitats. Second, using many noisy observations with a random sampling pattern is better than using fewer observations with a more structured sampling pattern. Third, using data from more engaged observers is generally better than less-engaged ones. Fourth, even though sociopolitical boundaries can have a significant effect on performance, training on diverse and extremely well-sampled high-resource regions like CA can still provide some predictive benefit for low-resource regions even if geographically far away.
@@
We first and foremost thank all of the participants who volunteered and contributed observations on iNaturalist. We further thank Noah Goodman and Gabriel Poesia for their comments and discussion, and we further thank CoCoLab for donating compute for this project. We also thank the TomKat Center for Sustainable Energy, The Fulbright Brasil Commission, Carnegie Science, the Howard Hughes Medical Institute, and the University of California, Berkeley for funding support for this research. This research was also funded by the NSF Graduate Research Fellowship DGE-1656518 (L.G.), the TomKat Graduate Fellowship for Translational Research (L.G.), and the Krupp Internship Program for Stanford Students in Germany (E.S.). T.K. and S.S. acknowledge funding from the German Research Foundation (DFG) within the framework of BigPlantSens (Assessing the Synergies of Big Data and Deep Learning for the Remote Sensing of Plant Species; project no. 444524904) and PANOPS (Revealing Earth’s plant functional diversity with citizen science; project no. 504978936). This research was further funded by the Office of the Director of the National Institutes of Health’s Early Investigator Award (1DP5OD029506-01, M.E.-A.). Compute for this project was performed on the Calc cluster at Carnegie Science and the Stanford SC Compute Cluster.